OpenStudy (anonymous):

Find the value of p that makes the linear graph y=p-3x pass through the point where the lines 4x-y=6 and 2x-5y=12 intersect.

OpenStudy (anonymous):

@Psymon

OpenStudy (anonymous):

@Compassionate

OpenStudy (***[isuru]***):

Hi, first make 2 equations from 4x - y = 6 and 2x -5y =12... 4x - y = 6 ------(1) 2x - 5y = 12 -----(2) Now .. solve them taking them as a system of equation ... and the values u get by solving them r coordinates of the intersection point of those 2 lines.. (2) x 2 4x -10y =24 -----(3) (3) - (1), 4x -y - [ 4x -10y ] = 18 9y = 18 y = 2 and substitute y = 2 in (1), 4x - y = 6 4x = 6+y 4x = 8 x = 2 So ... the intersection point of those 2 lines r (2,2) now if the function y = p-3x passes through that lpoin then the co-ordinates (2,2) should satisfy the function.. which mean.. 2 = p - (3x2) p = 8 yep !! that's it.. Hope this will help ya!!!

OpenStudy (***[isuru]***):

Did ya get it XD ?

OpenStudy (anonymous):

wait let me process it all in lol:p

OpenStudy (***[isuru]***):

sure...ya can take ur time !!lol

OpenStudy (anonymous):

I got y=-2 in the points

OpenStudy (anonymous):

@***[ISURU]***